Question 7. Solution needs to be handed in A given company has the following projected FCFs: Year FCF (Millions-SEK) 21 32 21 23 Suppose that the weighted average cost of capital is equal to 7% and that the free cash flows will grow at a rate equal to 25% beyond year 4. Suppose further that its debt is equal to 250 Million SEK and the number of shares is 30 million. What is the share price? (Round your answer to two decimal digits)

hi please can soneone explain how to solve this step by step
Answer #1

FCF1 = 21
FCF2 = 32
FCF3 = 21
FCF4 = 23

Growth rate, g = 2.50%
WACC = 7%

FCF5 = FCF4 * (1 + g)
FCF5 = 23 * 1.025
FCF5 = 23.575

Horizon Value, V4 = FCF5 / (WACC - g)
Horizon Value, V4 = 23.575 / (0.07 - 0.025)
Horizon Value, V4 = 523.889

Value of Firm, V0 = 21/1.07 + 32/1.07^2 + 21/1.07^3 + 23/1.07^4 + 523.889/1.07^4
Value of Firm, V0 = 481.94

Value of Firm = 481.94 Million SEK

Value of Equity = Value of Firm - Value of Debt
Value of Equity = 481.94 Million SEK - 250 Million SEK
Value of Equity = 231.94 Million SEK

Share Price = Value of Equity / Number of shares
Share Price = 231.94 Million SEK / 30 Million
Share Price = 7.73 SEK